OpenStudy (anonymous):

g(x) = -x² + 1 How do you find the Domain, Range, y-intercept and x-intercept?

OpenStudy (triciaal):

domain the values of x that will give a y value and therefore satisfy g(x) range the y values included the lowest value to the highest number included that satisfies the function y-intercept the value when x = 0 x-intercept the value when y= 0

OpenStudy (triciaal):

g(x) = -x² + 1 g(0) = -(0)^2 + 1 = 1 point (0,1) when g(x) = 0 then -x^2 + 1 = 0 1-x^2 = 0 x = +/- 1 points (-1, 0) and (1, 0)

OpenStudy (triciaal):